OpenStudy (anonymous):

PLEASE HELP :( PLEASEEE :/ Which of the following points are on the line given by the equation y = -3x +5? Check all that apply. A. (-2, 11) B. (3, -14) C. (2, -1) D. (-3, 14)

OpenStudy (anonymous):

im a little confused :/

OpenStudy (anonymous):

use the quadractic formula\[y=mx+b\] start with the point on the y- axis (the b in the equation) then use the rise and run of the m.. on graph.. use should be able to see it. If you give me what u think it might be, I'll let you know if you're right

OpenStudy (anonymous):

That's the slope y=mx+b

OpenStudy (anonymous):

Let me give you a little hint.

OpenStudy (anonymous):

b is the point on your y-axis

OpenStudy (anonymous):

So one coordinate is (0, 5)

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Wow. Only one doesn't apply o:

OpenStudy (anonymous):

Don't give it to her

OpenStudy (anonymous):

Oreo do you understand?

OpenStudy (mathstudent55):

Take the x- and y-coordinates of each point and plug them in for x and y in the equation of the line. If that makes the equation is true, then the point is on the line.

OpenStudy (anonymous):

I'm not, i'm waiting for everyone to say what their going to say. Lol.

OpenStudy (anonymous):

oops that's right.. so sorry.. quadractic is \[a^{2}+b^{2}=c ^{2}\] the explaination on how to use the slope is correct

OpenStudy (anonymous):

@mathstudent55 Good Explanation :3

OpenStudy (anonymous):

@Oreoo are you there?????

OpenStudy (anonymous):

How are we supposed to help

OpenStudy (anonymous):

Person waits til we say answer then says thanks. Usually how it goes.. Lol.

OpenStudy (mathstudent55):

@Mikeyy1992 Thanks

OpenStudy (anonymous):

im doing my work ok lol

OpenStudy (anonymous):

lol.. wish my question was getting this much traffic

OpenStudy (anonymous):

I don't see it I'll go help

OpenStudy (anonymous):

Mathstudent, gave you exactly what you needed to know to find the answer.

OpenStudy (anonymous):

Here, I'll show you a example.

OpenStudy (anonymous):

Say, you have (9, 20) as your choice. 9 is x, 20 is y. your equation is: y = -3x + 5 plug in: 20 = -3(9) + 5 20 = -22 Doesn't equal; means it's not on the line.

OpenStudy (anonymous):

Now, i'll show you one that applies, but the rest you will have to plug in on your own and ill confirm if you did them correctly. A) -2, 11 -2 is x 11 is y y = -3x + 5 11 = -3(-2) + 5 11 = 6 + 5 11 = 11 Equal amount; is a point on the line according with the equation given. Now try the 3 other choices and tell me what you get.

OpenStudy (anonymous):

is a?

OpenStudy (anonymous):

i mean i got a

OpenStudy (anonymous):

Hmm, you must read your question again. CHOOSE ALL THAT APPLY! So, you must check the other 3.

OpenStudy (anonymous):

11=-3(-2)+5 11=11

OpenStudy (anonymous):

oh yeaaa lol i forgot

OpenStudy (anonymous):

Now, try the other 3 and tell me what you get.

OpenStudy (anonymous):

i got all except for b

OpenStudy (anonymous):

Perfect.

OpenStudy (anonymous):

Medal, please?c:

OpenStudy (anonymous):

Thanks && suree

OpenStudy (anonymous):

Welcome. Have a good day or night!